OpenStudy (anonymous):

Solve for x in the following proportion. Round answer to nearest tenth. 27 : 10=x: 67

jhonyy9 (jhonyy9):

27/10 =x/67 --- 2,7 =x/67 --- multiply both sides with 67 and you get (67)(2,7)=x x=180,9